OpenStudy (anonymous):

Y=2x+3 2y=4x+6 the system of equations has __ solution no one infinity

OpenStudy (anonymous):

y/2y = 1/2 2x/4x = 1/2 & 3/6 = 1/2 from above it is obvious that the ratio of the coefficients of same kind of variables and constants are equal therefor given system of equations has infinite (many) solutions.

OpenStudy (anonymous):

@Lesliem